OpenStudy (anonymous):

Just check, plz! Jenna does 60 question in five days. On each day, she completes 3 more problems than the day before. How many problems did she do each day? Day 1___ Day 2____ Day 3___ ... Day 5___ So here is what I got: Fist day: 48 Second day:51 Third Day: 54 Fourth day:57 Ffith day: 60------} Is that correct?

OpenStudy (anonymous):

I think this means that she does a total of 60 questions over five days. So on the fifth day she is doing a + 12 questions (when a is the number of questions done on the first day), on the fourth day she is doing a + 9, third day is a + 6, second day is a + 3, and the first day is a. a + a + 3 + a + 6 + a + 9 + a + 12 = 60 combine like terms 5a + 30 = 60 5a = 30 a = 6 questions

OpenStudy (anonymous):

so on day 1, she does 6 questions?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

oky doky

jimthompson5910 (jim_thompson5910):

Yes echec is correct. Day 1: 6 problems Day 2: 6+3 = 9 problems Day 3: 9+3 = 12 problems Day 4: 12+3 = 15 problems Day 5: 15+3 = 18 problems Add up the individual counts: 6+9+12+15+18 = 60 So this checks the answer.

OpenStudy (anonymous):

thank you